Please Help!
Tom borrowed $2,000 from his father and agreed to pay a simple interest rate of 5.5%. After some time had passed, he paid his father $2,302.50. How long did it take Tom to pay back the loan, including interest?
Thank you for your time!

Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

he borrowed 2000.
he agreed to pay a simple interest rate of 5.5% per year.
he paid his father 2302.5 after some time.

with simple interest, the formula is:

f = p * r * n + p

f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

you have:
f = 2302.5
p = 2000
r = .055 per year
you want to solve for n.

formula becomes:

2302.5 = 2000 * .055 * n + 2000

subtract 2000 from both sides of this equation to get:

302.5 = 2000 * .055 * n

simplify to get:

302.5 = 110 * n

divide both sides of this equation by 110 to get:

302.5/110 = n

solve for n to get:

n = 2.75 years.

it took tom 2.75 years to pay back the loan.

go back to your original equation of:

2302.5 = 2000 * .055 * n + 2000

replace n with 2.75 to get:

2302.5 = 2000 * .055 * 2.75 + 2000

evaluate the equation to get:

2302.5 = 2302.5

solution looks good.

it took tom 2.75 years to pay back the loan.





Answer by MathTherapy(10552)


Please Help!
Tom borrowed $2,000 from his father and agreed to pay a simple interest rate of 5.5%. After some time had passed, he paid his father $2,302.50. How long did it take Tom to pay back the loan, including interest?
Thank you for your time!
Use the formula, A = P(1 + TR), where:
A = Accumulated loan amount/future value, or amount repaid at end of term ($2,302.50, in this case)
P = Principal, or original amount borrowed ($2,000, in this case)
T = Time, in years (Unknown, in this case)
R = Annual interest rate (5.5%, or .055, in this case)
Solve for T, the time, in years.